OpenStudy (anonymous):

2(x + 6) − (x + 1) = 16

OpenStudy (anonymous):

your equation can also be written like 2(x+6) + (-1)(x+1)= 16 so we start by distributing the 2 and distributing the -1 2x+12-x-1 = 16 can you solve this now?

OpenStudy (anonymous):

6-x-1=16 ? would that be right way to simplify ?

OpenStudy (anonymous):

no it would not rearrange it if you need to 2x+12-x-1 = 16 2x-x+12-1 = 16 (2x-x)+(12-1) = 16

OpenStudy (anonymous):

so is it 1x+11=16 ?

OpenStudy (anonymous):

then the answer is 5 right ?

OpenStudy (anonymous):

the entire answer would be "x=5" if you wrote just 5, then that wouldnt mean anything

OpenStudy (anonymous):

because you were solving for x

OpenStudy (anonymous):

obviously but you know what i mean right so no worries